Grade 11General Physics

Q.A train is moving on atrack at 30m/s.A ball is thrown from it perpendicular to direction of motion at 30m/s at 45degree from horizontal.Find the distance of ball from point of projection on train when it strikes the ground.
A)90m
B)90root3m
C)60m
D)60root3m

jiten
8 Years agoGrade 11

1 Answer

Khimraj
8 Years ago
time of flight will be 2*30*sin45/g = 3\sqrt{2}
distance = 30cos45*t = 30cos45*3\sqrt{2} = 90
hope it clears..............................